8. Let U be uniformly distributed on [0, L], where L has Exponential(A) distribution. Let V- L - U. What is the joint density function of U and V?
Answer #1

Given UIL (u and L~ Exp (A) . Also f_Lleft (l ight )=lambda e^{-lambda l};l>0 . The joint PDF

U,L (u,

Consider the transformation V=L-U,W=U. So U=W,L=W+V. The Jacobian of transformation is

1 0

The joint PDF of V,W is

/w. (v, u) = 1 × fvw (v, w)

The joint PDF of V,U is

f_{V,U}left (v,u ight )=left | J ight |f_{U,L}left (u,l ight ) f_{V,U}left (v,u ight )=1 imes rac{lambda e^{-lambda left ( u+v ight )}}{L} {color{Blue} f_{V,U}left (v,u ight )= rac{lambda e^{-lambda left ( u+v ight )}}{L};0<u<L,-L<v}